Operation Summary for March 23, 2011 - March 25, 2011 |
|
Wednesday March 23 The day shift began with Operations monitoring the Recycler stash (508.7E10), with NuMI off for target replacement, with the MiniBooNE and MTest off due to Transfer Hall and F-Sector accesses, and with the Tevatron off for D16-2 magnet replacement . During the day shift, Operations established beam to MiniBooNE at 11:10 AM and beam to MTest at 11:59 AM. MTest was approved for and began taking two spills of beam per minute at 3:00 PM. During the evening shift, Operations sent beam to MiniBooNE and MTest. Main Injector and Booster personnel performed studies throughout the shift.Thursday March 24 The midnight shift began with Operations monitoring the stash (493.9.9E10), and with MiniBooNE and MTest taking beam. During the midnight shift, Operations received alarms that indicated that the F0 expansion tank level was high and that the MRRF supply pressure was also too high. I:Q701 faulted off around the same time (1:00 AM) Operators investigated and found that water was coming out of the supply. Operators valved out the power supply and a decision was made to wait until morning to repair the power supply, this means no beam to MTest until repairs are completed. Beam was turned off at 6:00 AM for F-Sector, Booster and Main Injector accesses. During the day shift, F-Sector, Booster and Main Injector remained in access. During the evening shift, Operations established beam to MiniBooNE at 6:47 PM.Friday March 25 The midnight shift began with Operations sending beam to MiniBooNE and with the Main Injector, Linac, and Booster operating normally. During the midnight shift, Operations established beam to MTest at 4:39 AM.The Plans for Friday and the Weekend The plan for today and the weekend is to send beam to MiniBooNE and MTest. |
|
More Information Comments and Suggestions |
| last modified 03/25/2011 email Fermilab |
| Security, Privacy, Legal | |